# ProfOlaf: Semi-Automated Tool for Systematic Literature Reviews

### v1

@@ Abstract.
Systematic reviews and mapping studies are critical for synthesizing research, identifying gaps, and guiding future work, but they are often labor-intensive and time-consuming. Existing tools provide partial support for specific steps, leaving much of the process manual and error-prone. We present ProfOlaf, a semi-automated tool designed to streamline systematic reviews while maintaining methodological rigor. ProfOlaf supports iterative snowballing for article collection with human-in-the-loop filtering and uses large language models to assist in analyzing articles, extracting key topics, and answering queries about the content of papers. By combining automation with guided manual effort, ProfOlaf enhances the efficiency, quality, and reproducibility of systematic reviews across research fields. A video describing and demonstrating ProfOlaf is available at: https://youtu.be/4noUXfcmxsE
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrievalccs: Information systems Data management systemsccs: Computing methodologies Information extraction
@@ 1. Introduction
Despite their value, conducting systematic reviews is labor-intensive and time-consuming (Zhang and Babar, 2013; Van Dinter et al., 2021). The process typically requires conducting broad and comprehensive searches in multiple academic databases, followed by careful filtering of large volumes of articles against strict inclusion and exclusion criteria. After the collection is complete, the reviewers must also inspect and analyze the selected studies in depth, identify recurring research topics, and formulate answers to predefined research questions. Each of these steps demands sustained effort and precision.
Several methods and tools have been proposed to automate or support various steps of the systematic review process (Fabiano et al., 2024; Van Dinter et al., 2021). Existing solutions, such as reference managers or search engines, offer only partial assistance, leaving key tasks manual, time-consuming, and error-prone, reducing efficiency and reproducibility (Keele et al., 2007). More advanced approaches, such as AI-based tools, have shown promising results (Fabiano et al., 2024), but remain fragmented. Although these tools exist for individual steps of the process, no solution provides comprehensive support at all stages. Furthermore, achieving a sensible balance between manual effort and tool assistance is essential to ensure that precision is not compromised in the pursuit of reducing effort.
In response to these requirements, we propose ProfOlaf, a semi-automated tool designed to support and streamline the review process. ProfOlaf implements a structured methodology that adheres to established guidelines for systematic reviews. The methodology was designed with the goal of reducing human effort to the greatest extent possible. For the collection phase, ProfOlaf uses an iterative snowballing process to collect relevant articles, with each iteration involving human filtering supported by our tool. In the analysis phase, ProfOlaf integrates large language models (LLMs) to ease the analysis of the collected articles, enabling users to extract key topics from each article and query the model regarding its contents.
Figure 1. Overview of the ProfOlaf methodology. The search phase begins with an initial set of articles. Snowballing is applied, bibliographic information is retrieved, and venues may be optionally ranked. In the screening phase, article metadata is checked. Two or more human raters screen the articles by title and by full paper, with disagreements resolved collaboratively. This cycle continues until no new articles are identified. Duplicates are removed to form the final set of articles. In the data extraction, TopicGPT categorizes content into research topics, and LLM-based question answering is employed to extract structured insights. Manual inspection complements this step, ensuring a consolidated and reliable final set of extracted data.
ProfOlaf can assist researchers in conducting systematic reviews, thereby enhancing both the quality and the volume of such studies. ProfOlaf is open source and available at https://github.com/sr-lab/ProfOlaf, along with the experiments described in our evaluation.
@@ 2. ProfOlaf Overview
ProfOlaf is a Python tool with a structured and repeatable methodology for conducting systematic reviews, illustrated in Figure 1. The tool follows a methodology inspired by the guidelines proposed by Wohlin (Wohlin, 2014). The subsequent sections provide a detailed explanation of each step of the pipeline.
@@ 2.1. Initial Setup
To setup ProfOlaf  the user starts by defining a file containing an initial set of article titles. These articles are previously selected by the user and can be obtained through, for instance, a previous literature review on the topic being handled. The tool generates a database that stores all the metadata of the articles as well as the intermediate states of each article during the search.
@@ 2.2. Snowballing
After populating the database with the initial set of articles, ProfOlaf retrieves either the citations (forward snowballing), the references (backward snowballing), or both, for each article. The tool then compiles the bibliographic information for all the articles. Currently, ProfOlaf obtains article data from Google Scholar (Google, 2025), Semantic Scholar (Semantic Scholar, 2025), and DBLP (Schloss Dagstuhl - Leibniz Center for Informatics, 2025). However, the tool has been designed to ease the integration of additional search sources.
@@ 2.3. Metadata Screening
When the venue-ranking filter is applied, the user must first execute the step that assigns a ranking to all the venues present in the collected articles. To help the user with this manual task, ProfOlaf first uses cosine similarity to find venues previously ranked that are similar to the one being classified. Then, the tool searches for the venue in venue ranking databases and presents the top results for each database. For each result, ProfOlaf outputs the title of the venue, its registered ranking, and the cosine similarity score of the search. Currently, ProfOlaf searches both in the Scimago webpage and in a local CORE ranking table, but this functionality is easily extendable to other venue ranking databases.
2.4. Manual Screening
@@
As such, in this phase, the user is prompted to further refine the results, starting with removing irrelevant articles by title and then by the full content of the article. For this, ProfOlaf presents, for each article, its title as well as a url where the user can access it. After each step, the tool can be used to identify and display discrepancies between the user and other reviewers’ assessments. The reviewers must then discuss these disagreements and reach a consensus. The result of this discussion can then be inserted into the tool. With manual screening completed, the resulting set of articles is used as input to the snowballing phase (Section 2.2) to start a new iteration. If a given iteration does not produce new results, ProfOlaf consolidates the findings from all iterations into a single file. As a final precaution, the tool looks for articles with similar titles and checks with the user if they are in fact the same article under a different name. With this final check completed, the user obtains the final set of articles.
@@ 3. Evaluation
To evaluate ProfOlaf, we performed a small illustrative systematic literature review. We used the work of Ramos et al. (Ramos et al., [n. d.]) as our single seed article. We selected this paper because it won the Best Paper Award at the LLM4Code 2025 and because we are familiar with the topic of the paper.
@@ 3.1. Search and Screening Phase
The results of all iterations are summarized in Table 1. We calculate the efficiency measure for systematic literature reviews, a metric used by Wohlin (Wohlin, 2014) in their work to evaluate the amount of noise in the search. This metric is the number of included articles relative to the total number of candidate articles examined. The efficiency by iteration and the final efficiency are both represented in Table 1. As shown in Table 1, early iterations yielded a relatively higher proportion of relevant studies, while later iterations were characterized by a substantial increase in rejections. This pattern illustrates the phenomenon of diminishing returns in the search expansion process.
Table 1. Paper retrieval and screening results across iterations
[…]
@@
7 19 19 + 0 0 0
Total 1009 823 + 76 110 0.11
After screening, we obtained 111 articles, including the seed article by Ramos et al. (Ramos et al., [n. d.]). The final set of articles consisted of 108 articles, after the removal of three duplicates.
@@ 3.2. Data Extraction Phase
We selected two tasks for the Topic Modeling module: identifying the programming languages explored and the topics studied; and one task for the Task Assistant module, article summarization. We perform this evaluation with gpt-5-nano.
@@ 3.2.1. Topics Studied
Our topic modeling module identified 17 topics. To evaluate the capability of TopicGPT in assigning correct topics to articles, we constructed a ground truth by manually labeling each article. The first and second authors received the list of topics generated and were asked to assign topics to each article. Subsequently, the raters met to resolve disagreements and establish the final list of topic assignments. The process took in total 6.45 hours. The ground truth was compared with the TopicGPT results, which achieved a precision of 0.547 and a recall of 0.650.
The raters noted several observations. The topics were generally accurate, as it was possible to assign at least one topic to the majority of articles. However, some important topics were absent from the generated set, including Benchmark for Code Generation/Repair, Education, and Code Translation. When comparing the module’s assignments with the manual labels, we observed that the model frequently assigned both Code Generation and Automated Program Repair, whereas the raters selected only Automated Program Repair, treating it as a more specific instance of the broader category. Such cases likely contributed to the observed low precision score.
@@ 3.2.2. Programming Languages Used
For this task, the topic modeling module identified 40 programming languages. As in the previous task, assignments were performed both automatically by the tool and manually by the same raters. However, in this case, the raters were not restricted to the module’s generated set of answers. The entire process required 3.88 hours. The module achieved an average precision of 0.590 and an average recall of 0.710.
@@ 3.2.3. Task Assistant
Both raters generally agreed the summaries were accurate and free from hallucinations. Coverage was the lowest-scoring criterion, as some summaries did not fully capture the most important details; however, this limitation was minor (average score of 4.333). The other criteria received high scores, suggesting that the summaries were typically well-structured, easy to follow, and conveyed key points with minimal redundancy. Coherence/Structure had the highest standard deviation, indicating that the model’s ability to organize ideas varied the most across summaries.
Table 2. Evaluation of summary quality across 4 different criteria.
[…]
@@ 3.2.4. Discussion
The results indicate that while LLMs can deliver satisfactory outcomes for tasks such as summarization, more advanced models and methodologies are needed before they can be fully trusted for tasks such as topic modeling. For such cases, we argue that LLMs are best employed in a human-in-the-loop setting, either as assistants that facilitate manual work or as tools subject to human critique. For example, instead of manually identifying topics from scratch, a human could validate and refine the topic list generated by TopicGPT, as well as the assignments it produces. This approach substantially reduces effort while maintaining reliability.
@@ 5. Conclusion
ProfOlaf addresses key challenges in systematic reviews by combining iterative snowballing with LLM-assisted analysis, striking a balance between automation and human oversight. By improving efficiency, rigor, and reproducibility, it enables researchers to conduct higher-quality reviews with reduced effort. Given the rapidly growing volume of research in software engineering, ProfOlaf can be valuable in helping SE researchers keep pace with the field.

### v2

@@ Abstract.
Systematic reviews and mapping studies are critical to synthesize research, identify gaps, and guide future work, but are often labor-intensive and time-consuming. Existing tools provide partial support for specific steps, leaving much of the process manual and error-prone. We present ProfOlaf, a semi-automated tool designed to streamline systematic reviews while maintaining methodological rigor. ProfOlaf supports iterative snowballing for article collection with human-in-the-loop filtering and uses large language models to help select articles, extract key topics, and answer queries about the content of articles. By combining automation with guided manual effort, ProfOlaf enhances the efficiency, quality, and reproducibility of systematic reviews across research fields. ProfOlaf can be used both as a CLI tool and in web application format. A video demonstrating ProfOlaf is available at: https://youtu.be/R-gY4dJlN3s
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYbooktitle: Companion Proceedings of the 34th ACM Symposium on the Foundations of Software Engineering (FSE ’26), June 5–9, 2026, Montreal, Canadaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrievalccs: Information systems Data management systemsccs: Computing methodologies Information extraction
@@ 1. Introduction
Despite their value, conducting systematic reviews is labor-intensive and time-consuming (Zhang and Babar, 2013; Van Dinter et al., 2021). The process typically requires conducting broad and comprehensive searches in multiple academic databases, followed by careful filtering of large volumes of articles against predefined inclusion and exclusion criteria. After the collection is complete, the reviewers must also inspect and analyze the selected studies in depth, identify recurring research topics, and formulate answers to predefined research questions. Each of these steps demands sustained effort and precision.
Several methods and tools have been proposed to automate or support various steps of the systematic review process (Fabiano et al., 2024; Van Dinter et al., 2021). Existing solutions, such as reference managers or search engines, offer only partial assistance, leaving key tasks manual, time-consuming, and error-prone, reducing efficiency and reproducibility (Keele and others, 2007). More advanced approaches, such as AI-based tools, have shown promising results (Fabiano et al., 2024), but remain fragmented. Although these tools exist for individual steps of the process, no solution provides comprehensive support at all stages. Furthermore, achieving a sensible balance between manual effort and tool assistance is essential to ensure that precision is not compromised in the pursuit of reducing effort.
In response to these requirements, we propose ProfOlaf, a semi-automated tool designed to support and streamline the review process. ProfOlaf implements a structured methodology that adheres to established guidelines for systematic reviews. The methodology was designed with the goal of reducing human effort to the greatest extent possible. For the collection phase, ProfOlaf uses an iterative snowballing process to collect relevant articles, each iteration involving human-led filtering supported by our tool and optionally assisted by large language models (LLMs) as secondary reviewers. In the analysis phase, ProfOlaf integrates LLMs to ease the analysis of the collected articles, enabling users to extract key topics from each article and query the model regarding its contents.
Figure 1. Overview of the ProfOlaf methodology.
ProfOlaf assists researchers in conducting systematic reviews, improving both their quality and scalability. It is open source and available at https://github.com/sr-lab/ProfOlaf or at https://doi.org/10.5281/zenodo.18386217 (as a container), together with the experimental artifacts used in our evaluation. ProfOlaf is provided as both a command-line interface and a web application.
@@ 2. ProfOlaf Overview
The ProfOlaf methodology is illustrated in Figure 1. This section provides a detailed explanation of each step of the pipeline.
@@ 2.1. Initial Setup
To setup a new search with ProfOlaf  the user starts by defining a file containing an initial set of article titles. These articles are previously selected by the user and can be obtained through, for instance, a previous literature review on the topic being handled. The tool receives this file and generates a database that stores all the metadata of the articles as well as the intermediate states of each article during the search.
@@ 2.2. Snowballing
After populating the database with the initial set of articles, ProfOlaf retrieves either the citations (forward snowballing), the references (backward snowballing), or both, for each article. The tool then compiles the bibliographic information for the entire set.
ProfOlaf obtains article data from Google Scholar (Google, 2025), Semantic Scholar (Semantic Scholar, 2025), and DBLP (Schloss Dagstuhl - Leibniz Center for Informatics, 2025). However, the tool has been designed to ease the integration of additional search sources. As multiple versions of the same paper may exist under slightly different titles, this stage also includes a mechanism to identify and present these versions, allowing the user to select which one to keep.
@@ 2.3. Metadata Screening
When the venue-ranking filter is applied, the user must first execute the step that assigns a ranking to all the venues present in the collected articles. To help the user with this manual task, ProfOlaf first uses cosine similarity to find venues previously ranked that are similar to the one being classified. Then, the tool searches for the venue in venue ranking databases and presents the top results for each database. For each result, ProfOlaf outputs the title of the venue, its registered ranking, and the cosine similarity score of the search. Currently, ProfOlaf searches both in the SCImago and CORE databases, but this functionality is easily extendable to other venue ranking databases.
2.4. Article Screening
@@
As such, in this phase, the user is prompted to further refine the results, starting with removing irrelevant articles by title and then by their full content. For this, ProfOlaf presents, for each article, its title as well as a url where the user can access it. After each step, the tool can be used to identify and display discrepancies between the user and other reviewers’ assessments. The reviewers must then discuss these disagreements and reach a consensus. The result of this discussion can then be inserted into the tool.
In addition, the tool offers the option of using an LLM as an auxiliary rater to provide additional input during the discussion. Section 3.2 presents an evaluation of this component.
With both screening phases completed, the resulting set of articles is used as input to the snowballing phase (Section 2.2) to start a new iteration. If an iteration does not produce new results, ProfOlaf consolidates the findings from all iterations into a single file.
@@ 3. Evaluation
To evaluate ProfOlaf, we conducted a small illustrative systematic review using Ramos et al. (Ramos et al., ) as the seed article. This paper was selected for its Best Paper Award at LLM4Code 2025 and our familiarity with its topic.
@@ 3.1. Search and Screening Phase
The results of all iterations are summarized in Table 1. We calculate the efficiency measure for systematic literature reviews, a metric used by Wohlin (Wohlin, 2014) to evaluate the amount of noise in the search. Efficiency is calculated as the number of included articles relative to the total number of candidate articles examined. The efficiency by iteration and the final efficiency are both represented in Table 1. After screening, we obtained 108 articles, including the seed article by Ramos et al. (Ramos et al., ).
3.2. Automated Screening
We used ProfOlaf, configured to use gpt5.2, to perform both screening stages using a concise topic description along with inclusion and exclusion criteria. To evaluate its performance, we compared the model’s decisions against the final consensus reached by the two original raters. We also introduced an additional independent rater (fifth author) to screen the articles, allowing us to assess how the LLM’s performance compared to that of an individual human. Table 2 presents these findings. The screening process encompassed 183 articles at the title level and 125 articles at the full-text level.
We can observe that the LLM had a performance comparable to that of a human, with slightly greater precision, but lower recall, which indicates that the LLM tended to be more conservative in its screening decisions, favoring correctness over coverage.
Table 1. Snowballing process results across iterations.
I Retrieved Rejected (Meta. + Screen.) Approved Efficiency
@@ 3.2. Automated Screening
7 19 19 + 0 0 0.00
Total 1009 823 + 76 108 0.11
  • Notes: I = iteration; Meta. = metadata; Screen. = Screening.
Table 2. Automated screening evaluation.
Rater Screening Accuracy Precision Recall F1 Score
Human Title 0.8361 0.8613 0.9147 0.8872
Full Content 0.8760 0.9314 0.9223 0.9268
LLM Title 0.8415 0.9098 0.8605 0.8845
Full Content 0.8800 0.9417 0.9151 0.9282
@@ 3.3. Data Extraction Phase
We selected two tasks for the Topic Modeling module: identifying the programming languages explored and the topics studied; and one task for the Task Assistant module, article summarization. In this evaluation, ProfOlaf is configured to use gpt-5.2.
@@ 3.3.1. Topics Studied
A set of ground-truth topics was defined by two raters and subsequently revised and assigned during the manual analysis of the article collection. This ground truth is used to evaluate the topic modeling module via two distinct analyses: topic generation and topic assignment.
Topic Generation
Our ground truth consisted of 22 topics. Starting from Program Repair as the initial seed, TopicGPT generated 19 topics in its first step, of which 12 matched the ground-truth topics (54%). Notably, two ground-truth topics (Code Performance and Green Software) were merged into a single topic (Software Performance and Energy Efficiency). In addition, one generated topic, Automated Test Case Generation, only partially corresponded to our label Specification Generation and was therefore excluded from the identified topics. During the refinement step, TopicGPT reduced the set to 14 topics and correctly identified 45% of the ground-truth topics, with the same two mismatches as before. The performance decline is attributable to TopicGPT’s removal of infrequent topics; consequently, software tasks that did not frequently appear, such as Code Translation, were discarded.
Topic Assignment
Comparing the module’s topic assignments with those of the human raters, we obtained a precision of 0.645 and a recall of 0.850.
An analysis of the topic assignment metrics reveals that the Benchmarks label accounted for the highest number of incorrect assignments, with 30 false positives. 47.6% of the predictions assigned to this label were incorrect. This behavior is explained by the frequent mention of benchmark suites in many papers, even when they do not represent a core contribution, which leads the model to over-assign this topic. Addressing these alone would result in an estimated precision improvement of approximately 9%.
In contrast, while the Human–AI Collaboration label produced fewer false positives in absolute terms, it exhibited the highest misidentification rate (76.9% of false positives). This trend is attributed to the model’s tendency to assign this label to articles that merely reference human comparisons or interactions, rather than focusing on substantive contributions or studies centered on Human–AI collaboration in software engineering.
@@ 3.3.2. Programming Languages Used
As before, during the analysis of the papers, the programming languages were manually identified. For this task, the topic modeling module identified 40 programming languages, while the raters identified 47. The module achieved an average precision of 0.590 and an average recall of 0.710.
@@ 3.3.3. Task Assistant
Both raters generally agreed the summaries were accurate and free from hallucinations. Coverage was the lowest-scoring criterion, as some summaries did not fully capture the most important details; however, this limitation was minor (average score of 4.333). The remaining criteria received high scores, suggesting that the summaries were typically well-structured, easy to follow, and conveyed key points with minimal redundancy. Coherence/Structure had the highest standard deviation, indicating that the model’s ability to organize ideas varied the most across summaries.
Table 3. Summary Quality Evaluation.
Criterion Mean Std Criterion Mean Std
@@ 3.3.4. Discussion
The results suggest that while human raters cannot be fully replaced during the screening process, employing an LLM as an additional rater can provide valuable supplementary insights, stimulate discussion, and support the refinement of inclusion and exclusion criteria during the snowballing phase.
For more complex tasks such as topic modeling, our findings indicate that current large language models (LLMs) and associated methodologies are not yet sufficiently reliable to operate autonomously. Instead, their effectiveness is maximized in a human-in-the-loop setting, where they function as assistive tools rather than independent decision-makers. In this framework, LLMs can generate an initial set of topics and corresponding assignments, which are then validated and refined by a human reviewer. This approach substantially reduces manual effort while maintaining methodological reliability. Notably, the models did not hallucinate and generated plausible topics, but these often diverged from the task definition. More task-aligned prompt design could improve topic quality and relevance and is left for future work.
In contrast, for less demanding tasks such as summarization, LLMs already demonstrate satisfactory performance, making them well suited for direct use with minimal human intervention, saving time during the article analysis process.
@@ 5. Conclusion
ProfOlaf addresses key challenges in systematic review, combining iterative snowballing with LLM-assisted screening and analysis, striking a balance between automation and human oversight. By improving efficiency, rigor, and reproducibility, it enables researchers to conduct higher-quality reviews with reduced effort. Given the rapidly growing volume of research in software engineering, ProfOlaf can be valuable in helping SE researchers keep pace with the field.